OpenStudy (melmadans12):

given:7y=8x-14;y=6 prove: x=7 two column proof:)

OpenStudy (melmadans12):

i get that part, but the rest is confusing :(

OpenStudy (rzlearner):

Ok, so it seems you're going to try to get X on one side to get the answer so you'd get rid of the Y (substitute it) and work out the math :)

OpenStudy (rzlearner):

Let me know if you can do the math, or I can go ahead and help you out a bit more. To be honest I got the answer from a quick google search so I have another link too lol

OpenStudy (melmadans12):

42=8x-14 42+14=8x-14+14 56=8x 56/8=8x/8 7=x, see like this makes no sense to me

OpenStudy (rzlearner):

Well you got the steps right and that is the right answer so are you asking HOW does that prove the answer is 7...instead of any other answer? -.-

OpenStudy (melmadans12):

no no, like how would I write that out into columns

OpenStudy (rzlearner):

ahhhh ok lol well for the written part I'd start with writing the given information so you could put something like: 1.) Y=6 Given 2)X=8 Given 3)42+14=8x-14+14 Isolate the X variable 4)etc 5)etc

OpenStudy (rzlearner):

Just explain what you did for each step

OpenStudy (melmadans12):

7y=8x-14;y=6 = given is that right?

OpenStudy (melmadans12):

?

OpenStudy (rzlearner):

Yeah its looks good, you're just stating what you know for the first steps :)

OpenStudy (melmadans12):

can u help me with the rest of the column? I know it is asking a lot but this is for a test and I really need to bring my grade up

OpenStudy (melmadans12):

this is hopeless lol

OpenStudy (rzlearner):

You have a deadline or something? I'm at school right now so I am kind of multi tasking here lol

OpenStudy (melmadans12):

yeah it is due today and idk what im gonna do

OpenStudy (rzlearner):

You can do it!! Its not too tough, but yeah not knowing whether this its right or not gets annoying.

OpenStudy (rzlearner):

What time?

OpenStudy (melmadans12):

i have always been bad at math, and i really need a good grade

OpenStudy (melmadans12):

15 more mins

OpenStudy (rzlearner):

Crap...lol Well why don't you show me what you have and I can give you my input? Just make sure you all the steps you took to get to your answer and i'll help you write them out

OpenStudy (melmadans12):

As you got: 7y=8x-14 plug in the value of y=6 then : (7*6)=8x-14 42=8x-14 4214=8x 7=x

OpenStudy (melmadans12):

I*

likeabossssssss (likeabossssssss):

hey @melmadans12 \(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (melmadans12):

idk why i cant put a profile pic, but this is me:) lol in case u wax wonderin

OpenStudy (melmadans12):

thanks likeabosss lol

OpenStudy (melmadans12):

lmao ima fail math

OpenStudy (rzlearner):

1)Substitute the given value 6 for Y 2)Calculate both sides of the equation 3)Isolate the X variable of the equation by subtracting 14 from both sides 4)Solve for X

OpenStudy (melmadans12):

i usually dont ask for answers but at this point im desperate:/ thx rzlearner:)

OpenStudy (rzlearner):

welp thats what I got lol Very cute picture ^.^ And nooo don't say that haha you can do it!

OpenStudy (rzlearner):

I haven't been on in a while but happy to help :)

OpenStudy (melmadans12):

thanks:) and nooo i cant, ima just skip the questions lol

OpenStudy (rzlearner):

you're welcome, you'll be fine :) ha If its too hard then go ahead..its not worth the headache :) If you you need help with anything else just send a message. I might be on later? :)

OpenStudy (melmadans12):

ok sounds good,thanks